Write a system of equations for each situation.
a) Michael is three times older than his sister
Angela. In 1 year, Michael will be twice
as old as Angela. How old are the two
children today?
b) A $2 raffle ticket offers a bonus $1 early
bird draw. 400 tickets were sold for the
draw and a total of $894 was collected
from ticket sales. how many were
bought for $2 and how many were bought
for $3?

Answer by ankor@dixie-net.com(22740)

Write a system of equations for each situation.
a) Michael is three times older than his sister Angela.
m = 3a
In 1 year, Michael will be twice as old as Angela.
m + 1 = 3(a+1)
m + 1 = 3a + 3
m = 3a + 3 - 1
m = 3a + 2
Replace m with 3a, solve for a, then find m
How old are the two children today?
:
b) A $2 raffle ticket offers a bonus $1 early bird draw.
let x = no. of $2 tickets; let y = no. of $3 tickets
400 tickets were sold for the draw and
x + y = 400
y = (400-x)
a total of $894 was collected from ticket sales.
2x + 3y = 894
Replace y with (400-x), find x, then find y
how many were bought for $2 and how many were bought for $3?